Express the equation in logarithm form (use the same variable):
e^(3x + 25) = 60
my answer: ln(3x+25) = 60
therefore
x= 15
Is this right?

Answer by rothauserc(4718)

the natural logarithm form of
e^(3x + 25) = 60 is
3x + 25 = ln(60)
3x + 25 = 4.094344562
3x = −20.905655438
x = −6.968551813